OpenStudy (anonymous):

A motorcycle cost $12,000 when it was purchased. The value of a motorcycle decreases by 6% each year. Find the rate of decay each month and select the correct answer below.

OpenStudy (anonymous):

Wheres the answer choices

OpenStudy (anonymous):

I know you can use this formula A(T)=P+(1+r)^r/t or something like that lol

OpenStudy (michele_laino):

I'm sorry, I'm not good with financial mathematics

OpenStudy (anonymous):

:( Aw! It's fine:)

OpenStudy (anonymous):

Do you know anyone that can help?

OpenStudy (anonymous):

@welshfella @mathmate

OpenStudy (mathmate):

It depends on the definition of terms. If it depreciates 6% a year, most people will say that it depreciates 6/12=0.5% a month. However, it is not true, because 0.5% a month will end up with (1.005)^12=6.1678%. To get the real monthly depreciation, we need to solve the equation (1+i)^12=1.06 Take logs on both sides, and solve for i, we get i=e^(log(1.06)/12)-1=.004867550565343=0.486755% (I usually write log instead of ln. I write log10 when it's to the base 10). Check: (1.004867550565343)^12=1.0600000 so ok.